# Upho lattices II: ways of realizing a core

### v1

@@ 1. Introduction
Importantly, the core does not determine the upho lattice, in the sense that the same finite lattice can be the core of multiple upho lattices. For example, as depicted in Figure 1, the rank-two Boolean lattice $B_{2}$ is the core of two different upho lattices. [^1] We will show in Section 4 that these are the only two upho lattices which have $B_{2}$ as their core. In this paper, we study the question of in how many different ways a given finite lattice can be realized as a core. In a sense, we study whether an upho lattice can be represented by a finite amount of data.
@@ Acknowledgments
We thank the following people for useful comments related to this work: Ziyao Fu, Yibo Gao, Jon McCammond, Yulin Peng, and Yuchong Zhang. SageMath [12] was an important computational aid in this research.
@@ 2.1. Posets
Some posets $P$ we work with will be infinite. But every such $P$ will be at least locally finite, which means that all intervals $[x,y]\coloneqq\{z\in P\colon x\leq z\leq y\}$ in $P$ are finite. We recall that the Möbius function $\mu(x,y)$ of a locally finite poset $P$ can be defined recursively by $\mu(x,x)\coloneqq 1$ for all $x\in P$ and $\mu(x,y)\coloneqq-\sum_{x\leq z<y}\mu(x,z)$ for $x<y\in P$.
@@
For example, the Boolean lattice $B_{n}$ of subsets of $[n]\coloneqq\{1,\ldots,n\}$ ordered by inclusion is graded of rank $n$, and we have $F(B_{n};x)=(1+x)^{n}$ and $\chi^{*}(B_{n};x)=(1-x)^{n}$.
We use $\mathbb{N}\coloneqq\{0,1,2,\ldots\}$ for the natural numbers. An infinite poset $\mathcal{P}$ is $\mathbb{N}$-graded if it has a minimum $\hat{0}$ and it can be written as a disjoint union $\mathcal{P}=\bigsqcup_{i=0}^{\infty}\mathcal{P}_{i}$ such that every maximal chain in $\mathcal{P}$ is of the form $\hat{0}=x_{0}\lessdot x_{1}\lessdot x_{2}\lessdot\cdots$ with $x_{i}\in\mathcal{P}_{i}$. In this case, the poset’s rank function $\rho\colon\mathcal{P}\to\mathbb{N}$ is given by $\rho(x)=i$ if $x\in\mathcal{P}_{i}$. We say such a $\mathcal{P}$ is finite-type $\mathbb{N}$-graded if #$\#\mathcal{P}_{i}<\infty$ for all $i$. The rank generating and characteristic generating functions of such a $\mathcal{P}$ are then defined to be
$F(\mathcal{P};x)\coloneqq\sum_{p\in\mathcal{P}}x^{\rho(p)}\qquad\text{ and }\qquad\chi^{*}(\mathcal{P};x)\coloneqq\sum_{p\in\mathcal{P}}\mu(\hat{0},p)\,x^{\rho(p)}\,,$
@@ 3. Colorable upho posets, automorphisms, and finiteness
[…]
@@
So now suppose $w_{1},w_{2}\in S^{*}$ satisfy $w_{1}=w_{2}$ in $M$. Because the relations defining $M$ are homogeneous, $w_{1}$ and $w_{2}$ have the same length $n\geq 0$. We prove that $w_{1}=^{\prime}w_{2}$ by induction on $n$. If $n=0$ the claim is trivial, so suppose $n>0$. If for some $s\in S$ we have $w_{1}=su_{1}$ and $w_{2}=su_{2}$, then $u_{1}=u_{2}$ in $M$ because $M$ is left-cancellative; then by induction $u_{1}=^{\prime}u_{2}$, and so $w_{1}=^{\prime}w_{2}$. Otherwise, suppose without loss of generality that the word $w_{1}$ starts with the atom $s_{1}$, and $w_{2}$ starts with $s_{2}$. Let $x\coloneqq s_{1}\vee s_{2}\in\mathcal{L}$. In the proof of Lemma 3.1, we explain how elements of $M$ are identified with elements of $\mathcal{L}$. Let $u_{1},u_{2}\in S^{*}$ be words which when viewed as elements of $M$ are identified with $x\in\mathcal{L}$, and such that $u_{1}$ starts with $s_{1}$ while $u_{2}$ starts with $s_{2}$. Certainly $u_{1}=u_{2}$. But also, because the relations in $M^{\prime}$ equate words corresponding to saturated chains from two atoms to their join, we in fact have $u_{1}=^{\prime}u_{2}$. Now let $y\in\mathcal{L}$ be the element identified with $w_{1}=w_{2}\in M$. Since $y$ is an upper bound for $s_{1}$ and $s_{2}$, and $x$ is their least upper bound, we have that $x\leq y$. Thus, we can find a word $v\in S^{*}$ so that $u_{1}v=u_{2}v\in M$ is identified with $y\in\mathcal{L}$. Then $w_{1}=u_{1}v$, and both of these words start with $s_{1}$. Since $M$ is left-cancellative, we get an equality $\overline{w}_{1}=\overline{u}_{1}v$, where $\overline{w}_{1}$ is the result of removing the initial $s_{1}$ from $w_{1}$ and $\overline{u}_{1}$ is the result of removing the initial $s_{1}$ from $u_{1}$. But notice that the equality $\overline{w}_{1}=\overline{u}_{1}v$ involves words of length $n-1$, so by induction we have that $\overline{w}_{1}=^{\prime}\overline{u}_{1}v$, which then implies $w_{1}=^{\prime}u_{1}v$. A symmetric argument shows that $w_{2}=^{\prime}u_{2}v$. Finally, recalling $u_{1}=^{\prime}u_{2}$, we conclude $w_{1}=^{\prime}w_{2}$, as desired. ∎
@@
But how to construct upho colorings? Let $\mathcal{P}$ be an upho poset. A system of isomorphisms for $\mathcal{P}$ is a collection of isomorphisms $\varphi_{p}\colon V_{p}\to\mathcal{P}$ for each $p\in\mathcal{P}$. Note that, by definition of upper homogeneity, a system of isomorphisms for $\mathcal{P}$ exists. Let us say that a system of isomorphisms is compatible if
@@
Figure 2. Two different colored upho lattices with core $B_{2}$, as in 3.5 and 3.6. Also, compare Figure 1.
@@
Let $n\geq 1$ and let =$\mathcal{L}=\{\textrm{finite }\,S\subseteq\{1,2,\ldots\}\colon\max(S)<\#S+n\}$ (with the convention $\max(\varnothing)=0$), partially ordered by inclusion. Then $\mathcal{L}$ is another upho lattice with core $B_{n}$: see [8, Remark 8] and [9, §3.5.1]. A compatible system of isomorphisms for this $\mathcal{L}$ is $\varphi_{S}(T)=f_{S}(T\setminus S)$ where $f_{S}\colon\{1,2,\ldots\}\setminus S\to\{1,2,\ldots\}$ is the unique order-preserving bijection. And the corresponding monoid is then $M=\langle s_{1},\ldots,s_{n}\mid s_{i}s_{j-1}=s_{j}s_{i}\,\textrm{ for }\,1\leq i<j\leq n\rangle$. Compare this monoid to the one in 3.5. For $n=2$, the colored upho lattice $\mathcal{L}$ is depicted on the right in Figure 2.
@@
The upho lattices from 3.6 belong to a more general construction for any uniform sequence of supersolvable geometric lattices: see [9, §3]. These sequences include not just Boolean lattices, but also subspace lattices, partition lattices, etc. The conditions which go into the definition of “uniform sequence” imply the resulting systems of isomorphisms are compatible. Therefore, all the upho lattices in [9, §3] come from monoids. Also, because of semimodularity, all the monoids produced in this way will in fact have quadratic defining relations.
@@
If a finite graded lattice has no nontrivial automorphisms, then any upho lattice of which it is the core also has no nontrivial automorphisms.
@@
For many classes of finite, combinatorial structures, “most” members have no nontrivial automorphisms, in the sense that the proportion of such structures on $[n]$ with a nontrivial automorphism goes to 00 as $n\to\infty$. This is known to be true for finite graphs [5] and for finite posets [11], and we suspect it is true for finite lattices as well. Hence, Theorem 3.10 should apply to most lattices $L$. On the other hand, we also expect $\kappa(L)=0$ for most $L$. So it is reasonable to ask whether there are $L$ to which Theorem 3.10 applies but for which we also know that $\kappa(L)>0$. We provide an infinite sequence of such $L$, with both ranks and numbers of atoms going to infinity, in the following 3.12.
@@
We conclude this section by briefly explaining, in the following 3.13, how the techniques we developed here can sometimes be adapted to show that $\kappa(L)$ is finite for certain finite graded lattices $L$ that do have nontrivial automorphisms.
$a$$b$$c$$d$$e$
Figure 3. The lattice $L$ from 3.13 which has $\kappa(L)<\infty$ even though it has a nontrivial automorphism.
@@
Consider the rank-three lattice $L$ depicted in Figure 3. This $L$ has a nontrivial automorphism; it even has one which nontrivially permutes its atoms, as in 3.11. Nevertheless, it can be shown that $\kappa(L)$ is finite for this $L$. The basic idea is to consider what an upho lattice $\mathcal{L}$ that has $L$ as a core must look like up to rank three, and conclude that any such $\mathcal{L}$ must have no nontrivial automorphisms. Indeed, in any such upho lattice $\mathcal{L}$ with core $L$, there must be a unique element that covers the atom $b$ and that does not belong to the core $L$; call it $x$. By looking at the principal order filter generated by $b$, which must be isomorphic to $\mathcal{L}$ and so in particular have a copy of $L$ at the bottom, we see that for exactly one of the elements $d$ or $e$, the join of $x$ with this element has rank three. But this means that the way $\mathcal{L}$ looks up to rank three is asymmetric, and hence there are no automorphisms of $\mathcal{L}$ that nontrivially permute its atoms. By a similar argument as in the proof of Lemma 3.9, if $\mathcal{L}$ has no automorphisms that nontrivially permute its atoms, it has no nontrivial automorphisms at all. Hence, by Corollary 3.8 any such $\mathcal{L}$ is colorable, and therefore by Corollary 3.3 we indeed have $\kappa(L)<\infty$. We also believe $\kappa(L)>0$, because the monoid $M=\langle a,b,c\mid aa=bb,ba=ca\rangle$ should give an upho lattice with core $L$. This was mentioned in [9, Example 5.10] as well, although we must amend what is asserted there to merely state that $(M,\leq_{L})$ being an upho lattice with core $L$ is consistent with Conjecture 5.3 below.
@@
The first construction we call the dominating vertex construction of a lattice $\mathcal{D}_{n}$. We start by setting $P_{0}$ to be the one-element poset. Then, for each $i\geq 1$, we obtain $P_{i}$ from $P_{i-1}$ by doing the following:
  • we append a new element which covers all the elements of rank $i-1$;
  • then for each element $p$ of rank $i-1$, we also append $n-1$ additional new elements covering only $p$.
We call this the “dominating vertex construction” because of the element of rank $i$ which covers (“dominates”) all elements of rank $i-1$. It is clear that $P_{0}\subseteq P_{1}\subseteq\cdots$, with $P_{i}$ of rank $i$, so that we can define $\mathcal{D}_{n}\coloneqq\bigcup_{i=0}^{\infty}P_{i}$ to be the result of this construction. For example, the left side of Figure 4 depicts the first few ranks of $\mathcal{D}_{3}$.
The second construction we call the flip construction of a lattice $\mathcal{F}_{n}$. We start by setting $P_{0}$ to be the one-element poset and $P_{1}$ to be the “claw” poset with minimum $\hat{0}$, $n$ atoms, and no other elements. Then, for each $i\geq 2$, we obtain $P_{i}$ from $P_{i-1}$ by doing the following:
  • for each element $p$ of rank $i-2$, letting $q_{1},\ldots,q_{n}$ be the elements of rank $i-1$ covering $p$, we append a new element which covers exactly $q_{1},\ldots,q_{n}$;
  • then for each element $p$ of rank $i-1$, we also append enough additional new elements covering only $p$ to make $p$ be covered by exactly $n$ elements.
We call this the “flip construction” because the first step can be seen as taking the portion of the Hasse diagram between ranks $i-2$ and $i-1$ and placing a reflected copy above it. It is again clear that $P_{0}\subseteq P_{1}\subseteq\cdots$, with $P_{i}$ of rank $i$, so that we can define $\mathcal{F}_{n}\coloneqq\bigcup_{i=0}^{\infty}P_{i}$ to be the result of this construction. For example, the right side of Figure 4 depicts the first few ranks of $\mathcal{F}_{3}$.
@@
Let $\mathcal{L}$ be an upho lattice with core $M_{n}$ for some $n\geq 2$. If $\mathcal{L}$ agrees with $\mathcal{D}_{n}$ up to rank three, it must in fact be $\mathcal{D}_{n}$ (up to isomorphism). Ditto for $\mathcal{F}_{n}$.
@@
We first consider the dominating vertex construction. Fix an integer $k\geq 3$, and suppose that $\mathcal{L}$ is an upho lattice with core $M_{n}$ and that for each $i\leq k$, there is an element of rank $i$ in $\mathcal{L}$ that covers all elements of rank $i-1$ (i.e., $\mathcal{L}$ agrees with $\mathcal{D}_{n}$ up to rank $k$). Let $s_{1}$, …, $s_{n}$ be the atoms of $\mathcal{L}$. For any $1\leq i\leq n$, consider the order filter $V_{s_{i}}$ above $s_{i}$: it contains some set $S_{i}$ of vertices at rank $k$ in $\mathcal{L}$. These vertices are at rank $k-1$ in $V_{s_{i}}$, and since $V_{s_{i}}\cong\mathcal{L}$, we have by assumption that there is an element $x_{i}$ in $V_{s_{i}}$ that covers all vertices in $S_{i}$.
Let $u=s_{1}\vee\cdots\vee s_{n}$ be the element of rank $2$ in $\mathcal{L}$ that covers all the $s_{i}$. Consider the set $S$ of vertices at rank $k$ in $\mathcal{L}$ that are larger than $u$. Obviously, $S\subseteq S_{i}$ for all $1\leq i\leq n$, and hence every element in $S$ is covered by $x_{1},\ldots,x_{n}$. Since $k>2$ and $\mathcal{L}$ is upho and not a chain, #$\#S\geq 2$. That is, $S$ is a set of at least $2$ vertices in $\mathcal{L}$, all of which are covered by all of $x_{1},\ldots,x_{n}$. Since $\mathcal{L}$ is a lattice, this is only possible if $x_{1}=\cdots=x_{n}$, and this single element covers all of the vertices in all of the $S_{i}$. Since every element at rank $k$ in $\mathcal{L}$ must be larger than some atom $s_{i}$, and therefore must belong to one of the sets $S_{i}$, it follows that the element $x_{1}$ covers all elements of rank $k$ in $\mathcal{L}$, so that $\mathcal{L}$ agrees with the $\mathcal{D}_{n}$ through rank $k+1$. The result now follows by induction.

### v2

@@ 1. Introduction
Importantly, the core does not determine the upho lattice, in the sense that the same finite lattice can be the core of multiple upho lattices. For example, as depicted in Figure 1, the rank-two Boolean lattice $B_{2}$ is the core of two different upho lattices. [^1] We will show in Section 4 that these are the only two upho lattices which have $B_{2}$ as their core. In this paper, we study the question of in how many different ways a given finite lattice can be realized as a core. In a sense, we study whether every upho lattice can be represented by a finite amount of data.
@@ Acknowledgments
We thank Ziyao Fu, Yibo Gao, Jon McCammond, Yulin Peng, David Speyer, and Yuchong Zhang for useful comments related to this work. SageMath [12] was an important computational aid in this research. Finally, we thank the anonymous referees for their careful reading of our paper, and useful comments.
@@ 2.1. Posets
Some posets $P$ we work with will be infinite. But every such $P$ will be at least locally finite, which means that all intervals $x,y]\coloneqq\{z\in P\colon x\leq z\leq y\}$ in $P$ are finite. We recall that the Möbius function $\mu(x,y)$ of a locally finite poset $P$ can be defined recursively by $\mu(x,x)\coloneqq 1$ for all $x\in P$ and $\mu(x,y)\coloneqq-\sum_{x\leq z<y}\mu(x,z)$ for $x<y\in P$.
@@
For example, the Boolean lattice $B_{n}$ of subsets of $n]\coloneqq\{1,\ldots,n\}$ ordered by inclusion is graded of rank $n$, and we have $F(B_{n};x)=(1+x)^{n}$ and $\chi^{*}(B_{n};x)=(1-x)^{n}$.
We use $\mathbb{N}\coloneqq\{0,1,2,\ldots\}$ for the natural numbers. An infinite poset $\mathcal{P}$ is $\mathbb{N}$-graded if it has a minimum $\hat{0}$ and it can be written as a disjoint union $\mathcal{P}=\bigsqcup_{i=0}^{\infty}\mathcal{P}_{i}$ such that every maximal chain in $\mathcal{P}$ is of the form $\hat{0}=x_{0}\lessdot x_{1}\lessdot x_{2}\lessdot\cdots$ with $x_{i}\in\mathcal{P}_{i}$. In this case, the poset’s rank function $\rho\colon\mathcal{P}\to\mathbb{N}$ is given by $\rho(x)=i$ if $x\in\mathcal{P}_{i}$. We say such a $\mathcal{P}$ is finite-type $\mathbb{N}$-graded if $\mathcal{P}_{i}<\infty$ for all $i$. The rank generating and characteristic generating functions of such a $\mathcal{P}$ are then defined to be
[…]
@@ 3. Colorable upho posets, automorphisms, and finiteness
$M\coloneqq\left\langle s_{1},\ldots,s_{r}\mid\parbox{234.87749pt}{\begin{center} $c(\hat{0}=x_{0}\lessdot x_{1})c(x_{1}\lessdot x_{2})\cdots c(x_{k-1}\lessdot x_{k}=p)=$ \\ $c(\hat{0}=y_{0}\lessdot y_{1})c(y_{1}\lessdot y_{2})\cdots c(y_{k-1}\lessdot y_{k}=p)$ \end{center}}\right\rangle$
@@
Remark 3.2.
From now on, we will frequently use the correspondence between the elements of a colored upho poset $\mathcal{P}$ and the elements of its associated monoid $M$: the element $p\in\mathcal{P}$ in the poset corresponds to the element $c(x_{0}\lessdot x_{1})\cdots c(x_{k-1}\lessdot x_{k})\in M$ in the monoid, where $\hat{0}=x_{0}\lessdot x_{1}\lessdot\cdots\lessdot x_{k}=p$ is any maximal chain from $\hat{0}$ to $p$.
@@
So now suppose $w_{1},w_{2}\in S^{*}$ satisfy $w_{1}=w_{2}$ in $M$. Because the relations defining $M$ are homogeneous, $w_{1}$ and $w_{2}$ have the same length $n\geq 0$. We prove that $w_{1}=^{\prime}w_{2}$ by induction on $n$. If $n=0$ the claim is trivial, so suppose $n>0$. If for some $s\in S$ we have $w_{1}=su_{1}$ and $w_{2}=su_{2}$, then $u_{1}=u_{2}$ in $M$ because $M$ is left-cancellative; then by induction $u_{1}=^{\prime}u_{2}$, and so $w_{1}=^{\prime}w_{2}$. Otherwise, suppose without loss of generality that the word $w_{1}$ starts with the atom $s_{1}$, and $w_{2}$ starts with $s_{2}$. Let $x\coloneqq s_{1}\vee s_{2}\in\mathcal{L}$. Let $u_{1},u_{2}\in S^{*}$ be words which when viewed as elements of $M$ are identified with $x\in\mathcal{L}$ (as in Remark 3.2), and are such that $u_{1}$ starts with $s_{1}$ while $u_{2}$ starts with $s_{2}$. Certainly $u_{1}=u_{2}$. But also, because the relations in $M^{\prime}$ equate words corresponding to saturated chains from two atoms to their join, we in fact have $u_{1}=^{\prime}u_{2}$. Now let $y\in\mathcal{L}$ be the element identified with $w_{1}=w_{2}\in M$. Since $y$ is an upper bound for $s_{1}$ and $s_{2}$, and $x$ is their least upper bound, we have that $x\leq y$. Thus, we can find a word $v\in S^{*}$ so that $u_{1}v=u_{2}v\in M$ is identified with $y\in\mathcal{L}$. Then $w_{1}=u_{1}v$, and both of these words start with $s_{1}$. Since $M$ is left-cancellative, we get an equality $\overline{w}_{1}=\overline{u}_{1}v$, where $\overline{w}_{1}$ is the result of removing the initial $s_{1}$ from $w_{1}$ and $\overline{u}_{1}$ is the result of removing the initial $s_{1}$ from $u_{1}$. But notice that the equality $\overline{w}_{1}=\overline{u}_{1}v$ involves words of length n1n-1, so by induction we have that $\overline{w}_{1}=^{\prime}\overline{u}_{1}v$, which then implies $w_{1}=^{\prime}u_{1}v$. A symmetric argument shows that $w_{2}=^{\prime}u_{2}v$. Finally, recalling $u_{1}=^{\prime}u_{2}$, we conclude $w_{1}=^{\prime}w_{2}$, as desired. ∎
@@
But how to construct upho colorings? Let $\mathcal{P}$ be an upho poset. A system of isomorphisms for $\mathcal{P}$ is a collection of isomorphisms $\varphi_{p}\colon V_{p}\to\mathcal{P}$ for each $p\in\mathcal{P}$. (Such systems always exist, by the definition of upper homogeneity.) Let us say that a system of isomorphisms is compatible if
@@
Figure 2. Two different colored upho lattices with core $B_{2}$, as in 3.6 and 3.7. Also, compare Figure 1.
@@
Let $n\geq 1$ and let $\mathcal{L}=\{\textrm{finite }\,S\subseteq\{1,2,\ldots\}\colon\max(S)<\#S+n\}$ (with the convention $\max(\varnothing)=0$), partially ordered by inclusion. Then $\mathcal{L}$ is another upho lattice with core $B_{n}$: see [8, Remark 8] and [9, §3.5.1]. A compatible system of isomorphisms for this $\mathcal{L}$ is $\varphi_{S}(T)=f_{S}(T\setminus S)$ where $f_{S}\colon\{1,2,\ldots\}\setminus S\to\{1,2,\ldots\}$ is the unique order-preserving bijection. And the corresponding monoid is then $M=\langle s_{1},\ldots,s_{n}\mid s_{i}s_{j-1}=s_{j}s_{i}\,\textrm{ for }\,1\leq i<j\leq n\rangle$. Compare this monoid to the one in 3.6. For $n=2$, the colored upho lattice $\mathcal{L}$ is depicted on the right in Figure 2.
@@
The upho lattices from 3.7 belong to a more general construction for any uniform sequence of supersolvable geometric lattices: see [9, §3]. These sequences include not just Boolean lattices, but also subspace lattices, partition lattices, etc. The conditions which go into the definition of “uniform sequence” imply the resulting systems of isomorphisms are compatible. Therefore, all the upho lattices in [9, §3] come from monoids. Also, because of semimodularity, all the monoids produced in this way will in fact have quadratic defining relations.
@@
If a finite graded lattice $L$ has no nontrivial automorphisms, then any upho lattice with core $L$ also has no nontrivial automorphisms.
@@
For many classes of finite, combinatorial structures, “most” members have no nontrivial automorphisms, in the sense that the proportion of such structures on $[n]$ with a nontrivial automorphism goes to 0 as $n\to\infty$. This is known to be true for finite graphs [5] and for finite posets [11], and we suspect it is true for finite lattices as well. Hence, Theorem 3.11 should apply to most lattices $L$. On the other hand, we also expect $\kappa(L)=0$ for most $L$. So it is reasonable to ask whether there are $L$ to which Theorem 3.11 applies but for which we also know that $\kappa(L)>0$. We provide an infinite sequence of such $L$, with both ranks and numbers of atoms going to infinity, in the following 3.13.
@@
We conclude this section by briefly explaining, in the following 3.14, how the techniques we developed here can sometimes be adapted to show that $\kappa(L)$ is finite for certain finite graded lattices $L$ that do have nontrivial automorphisms.
aabbccddee
Figure 3. The lattice $L$ from 3.14 which has $\kappa(L)<\infty$ even though it has a nontrivial automorphism.
@@
Consider the rank-three lattice $L$ depicted in Figure 3. This $L$ has a nontrivial automorphism; it even has one which nontrivially permutes its atoms, as in 3.12. Nevertheless, it can be shown that $\kappa(L)$ is finite for this $L$. The basic idea is to consider what an upho lattice $\mathcal{L}$ that has $L$ as a core must look like up to rank three, and conclude that any such $\mathcal{L}$ must have no nontrivial automorphisms. Indeed, in any such upho lattice $\mathcal{L}$ with core $L$, there must be a unique element that covers the atom $b$ and that does not belong to the core $L$; call it $x$. By looking at the principal order filter generated by $b$, which must be isomorphic to $\mathcal{L}$ and so in particular have a copy of $L$ at the bottom, we see that for exactly one of the elements $d$ or $e$, the join of $x$ with this element has rank three. But this means that the way $\mathcal{L}$ looks up to rank three is asymmetric, and hence there are no automorphisms of $\mathcal{L}$ that nontrivially permute its atoms. By a similar argument as in the proof of Lemma 3.10, if $\mathcal{L}$ has no automorphisms that nontrivially permute its atoms, it has no nontrivial automorphisms at all. Hence, by Corollary 3.9 any such $\mathcal{L}$ is colorable, and therefore by Corollary 3.4 we indeed have $\kappa(L)<\infty$. We also believe $\kappa(L)>0$, because the monoid $M=\langle a,b,c\mid aa=bb,ba=ca\rangle$ should give an upho lattice with core $L$, conditional on the validity of Conjecture 5.3 below in this case. [^3] This example was considered in [9, Example 5.10], where the conjecture was implicitly assumed.
@@
The first construction we call the dominating vertex construction, which produces a lattice we denote $\mathcal{D}_{n}$. We start by setting $P_{0}$ to be the one-element poset. Then, for each $i\geq 1$, we obtain $P_{i}$ from $P_{i-1}$ by doing the following:
  • we append a new element which covers all the elements of rank i1i-1;
  • then for each element $p$ of rank i1i-1, we also append n1n-1 additional new elements covering only $p$.
We call this the “dominating vertex construction” because of the element of rank $i$ which covers (“dominates”) all elements of rank i1i-1. It is clear that $P_{0}\subseteq P_{1}\subseteq\cdots$, with $P_{i}$ of rank $i$, so that we can define $\mathcal{D}_{n}\coloneqq\bigcup_{i=0}^{\infty}P_{i}$ to be the result of this construction. For example, the left side of Figure 4 depicts the first few ranks of $\mathcal{D}_{3}$.
The second construction we call the flip construction, which produces a lattice we denote $\mathcal{F}_{n}$. We start by setting $P_{0}$ to be the one-element poset and $P_{1}$ to be the “claw” poset with minimum $\hat{0}$, $n$ atoms, and no other elements. Then, for each $i\geq 2$, we obtain $P_{i}$ from $P_{i-1}$ by doing the following:
  • for each element $p$ of rank i2i-2, letting $q_{1},\ldots,q_{n}$ be the elements of rank i1i-1 covering $p$, we append a new element which covers exactly $q_{1},\ldots,q_{n}$;
  • then for each element $p$ of rank i1i-1, we also append enough additional new elements covering only $p$ to make $p$ be covered by exactly $n$ elements.
We call this the “flip construction” because the first step can be seen as taking the portion of the Hasse diagram between ranks i2i-2 and i1i-1 and placing a reflected copy above it. It is again clear that $P_{0}\subseteq P_{1}\subseteq\cdots$, with $P_{i}$ of rank $i$, so that we can define $\mathcal{F}_{n}\coloneqq\bigcup_{i=0}^{\infty}P_{i}$ to be the result of this construction. For example, the right side of Figure 4 depicts the first few ranks of $\mathcal{F}_{3}$.
@@
Let $\mathcal{L}$ be an upho lattice with core $M_{n}$ for some $n\geq 2$. If $\mathcal{L}$ agrees with $\mathcal{D}_{n}$ up to rank three, it must in fact be $\mathcal{D}_{n}$ (up to isomorphism). The same is true for $\mathcal{F}_{n}$.
@@
We first consider the dominating vertex construction. Fix an integer $k\geq 3$, and suppose that $\mathcal{L}$ is an upho lattice with core $M_{n}$ and that for each $i\leq k$, there is an element of rank $i$ in $\mathcal{L}$ that covers all elements of rank i1i-1. Note that this implies that $\mathcal{L}$ agrees with $\mathcal{D}_{n}$ up to rank $k$. Indeed, each element at rank $i$ other than the “dominating” one covers exactly one element of rank i1i-1: if another rank-$i$ element covered two or more elements of rank i1i-1, then these rank-(i1)(i-1) elements would have incomparable upper bounds at rank $i$; and each element of rank $i$ must cover an element of rank i1i-1 for the lattice to be graded. Moreover, each element at rank i1i-1 is covered by $n$ elements: one is the “dominating” element of rank $i$ and the other n1n-1 have no other lower covers.
Let $s_{1}$, …, $s_{n}$ be the atoms of $\mathcal{L}$. For any $1\leq i\leq n$, consider the order filter $V_{s_{i}}$ above $s_{i}$: it contains some set $S_{i}$ of vertices at rank $k$ in $\mathcal{L}$. These vertices are at rank k1k-1 in $V_{s_{i}}$, and since $V_{s_{i}}\cong\mathcal{L}$, we have by assumption that there is an element $x_{i}$ in $V_{s_{i}}$ that covers all vertices in $S_{i}$.
Let $u=s_{1}\vee\cdots\vee s_{n}$ be the element of rank $2$ in $\mathcal{L}$ that covers all the $s_{i}$. Consider the set $S$ of vertices at rank $k$ in $\mathcal{L}$ that are larger than $u$. Obviously, $S\subseteq S_{i}$ for all $1\leq i\leq n$, and hence every element in $S$ is covered by $x_{1},\ldots,x_{n}$. Since $k>2$ and $\mathcal{L}$ is upho and not a chain, $S\geq 2$. That is, $S$ is a set of at least $2$ vertices in $\mathcal{L}$, all of which are covered by all of $x_{1},\ldots,x_{n}$. Since $\mathcal{L}$ is a lattice, this is only possible if $x_{1}=\cdots=x_{n}$, and this single element covers all of the vertices in all of the $S_{i}$. Since every element at rank $k$ in $\mathcal{L}$ must be larger than some atom $s_{i}$, and therefore must belong to one of the sets $S_{i}$, it follows that the element $x_{1}$ covers all elements of rank $k$ in $\mathcal{L}$, so that $\mathcal{L}$ agrees with the $\mathcal{D}_{n}$ through rank $k+1$. The result now follows by induction.